Give your verdict. Present the final result in a simple language and underline the final

alpha=0.05.


[ Additional instructions: A) Only handwritten material on the answer sheet will be evaluated. De |

its screenshots. B) Write all the steps- including the hypothesis, and show calculations for co

answer sheet, even if Excel or any other software is used. If Excel or another software is used to

write the correct Excel function on the answer sheet.]



1
Expert's answer
2021-12-01T13:37:55-0500

for example:

Let’s say we need to determine if girls on average score higher than 600 in the exam. We have the information that the standard deviation for girls’ scores is 100. So, we collect the data of 20 girls by using random samples and record their marks. Finally, we also set our ⍺ value (significance level) to be 0.05.

In this example:

  • Mean Score for Girls is 641
  • The size of the sample is 20
  • The population mean is 600
  • Standard Deviation for Population is 100

"H_0:\mu=600"

"H_a:>600"


"t=\frac{\overline{x}-\mu}{\sigma/\sqrt n}=\frac{641-600}{100/\sqrt{20}}=1.83"

"df=n-1=19"

critical value:

"t_{crit}=1.729"


Since the "|t|>t_{crit}" , we can reject the null hypothesis and conclude based on our result that Girls on average scored higher than 600.